1.How long it will take R3000 Interest a 6% per annum simple interest to grow to R 4260
2.James borrows R1250 at 7.5% compound interest per annum for 3years. Calculate the total amounts he needs to repay.
3.If The exchange rate between us dollar And rand is R13,90= 1$ today.Calculate how much A pair of shoes will cost In rand value if it is marked 40$ in USA

Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52814)

.

One and ONLY ONE problem/question per post.



It is the RULE,  the  POLICY  and the  REQUIREMENT  of this forum.

It is written in this page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.


It is assumed that you read these rules before posting.

It is also assumed that you do understand what is written in that page and follow the rules.


Those who violate them,  work against their own interests.


//////////////////


Both the solution and the answer to problem 2 in post by @Theo are incorrect.



Answer by Theo(13342)

1.How long it will take R3000 Interest a 6% per annum simple interest to grow to R 4260?

sinple interest formula is f = p + p * r * n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time periods
formula becomes 4260 = 3000 + 3000 * .06 * n
subtract 3000 from both sides of the equation to get:
1260 = 3000 * .06 * n
divide both sides of the formula by (3000 * .06) to get:
1260 / (3000 * .06) = n
solve for n to get:
n = 7
confirm by replacing n with 7 in the original equation to get:
4260 = 3000 + 3000 * .06 * 7 which becomes:
4260 = 4260, confirming that n = 7 is correct.

2.James borrows R1250 at 7.5% compound interest per annum for 3years. Calculate the total amounts he needs to repay.

formula for this is f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period.
n is the number of time periods.
formula becomes:
f = 1250 * (1 + .075) ^ 3
simplify to get:
f = 1250 * 1.075 ^ 3
solve for f to get:
f = 1552.87, rounded to the nearest penny.
that's what he needs to repay at the end of the 3 loan period.

3.If The exchange rate between us dollar And rand is R13,90 = 1$ today.
Calculate how much A pair of shoes will cost In rand value if it is marked 40$ in USA

the exchange rate is 13.90 rand for every 1 us dollar.
if the shoes cost 40 us dollars, then the rand cost would be 13.90 * 40 = 556 rand.